OpenStudy (libanfarah):

There are 1644 students in central high school. There are 50 more girls than boys. How many boys are there

OpenStudy (schoolgeek15):

I got you hold on

OpenStudy (abb0t):

g = girls b = boys g + b = 1644 g = 50 + b (50 + b) + b = 1644 2b = 1594 b = 797 g + b = 1644 g + 797 = 1644 g = 847 847 girls, 797 boys.

OpenStudy (schoolgeek15):

2x+50=1644 that is the equaltion so do u know what to do now allright let me explain to u how i get the equaltion from 50 is the amount that boys more than girls and then u have to divide the 1644 by 2 because u have to know how boys and girls are there. here i am very bab at word explain i can explain to u in equation : first u have to do this : (1644-50) then u get 1594 so u have to divide 1594 by 2 to get the amount of boys and girls .because they have to be even now u should get 797 here 797 is the answer

likeabossssssss (likeabossssssss):

\(\Huge\color{#EB00FF}{\text{WELCOME}}\) \(\Huge\color{blue}{\text{TO}}\) \(\Huge\color{green}{\text{OPEN}}\)\(\Huge\color{purple }{\text{ STUDY!!!!!!!!!!!}}\) \(\Huge\heartsuit\)